OpenStudy (anonymous):

Please answer!!!!! Factor: 1) −5x3 + 10x2 2) 3x2 + 9x − 3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$−5x^3 + 10x^2$$

OpenStudy (anonymous):

I think that for 1 the GCF's are 5 and x^2?

OpenStudy (skullpatrol):

Good, you are right :)

OpenStudy (anonymous):

ok now would it be -5x^2(5+x^2)? Or something like that?

OpenStudy (anonymous):

I dont know the firt part before the parenthasis...

OpenStudy (anonymous):

parenthesis

OpenStudy (skullpatrol):

Multiply them using the distributive property: $$a(b + c) = ab + ac$$ where a is the first part before the parenthesis.

OpenStudy (anonymous):

uhh? :3

OpenStudy (anonymous):

−5x2(x − 2)? and 2 = 3(x2 + 3x − 1)?

OpenStudy (anonymous):

yeah watever thx...

OpenStudy (anonymous):

FOR (2) \[3x ^{2}+9x-3=3\left( x ^{2} +3x\right)-3=3\left( x ^{2}+3x+\frac{ 9 }{ 4 }-\frac{ 9 }{ 4 } \right)-3\] \[=3\left( x+\frac{ 3 }{ 2} \right)^{2}-\frac{ 27 }{4 }-3\] \[=3\left( x+\frac{ 3 }{ 2 } \right)^{2}-\frac{ 39 }{4 }\] \[=3[\left( x+\frac{ 3 }{ 2 } \right)^{2}-\frac{ 13 }{ 4 }]\] \[=3\left[ \left( x+\frac{ 3 }{ 2 } \right)^{2}-\left( \frac{ \sqrt{13} }{2 } \right)^{2}\right]\] \[=3\left( x+\frac{ 3 }{ 2 }+\frac{ \sqrt{13} }{2 } \right)\left(x+\frac{ 3 }{ 2 }-\frac{ \sqrt{13 } }{ 2} \right)\]